OpenStudy (anonymous):

which value of x gives you sin(3(x+120)°) = 1/2

OpenStudy (anonymous):

oops sorry a)170 b)175 c)180 d)185

OpenStudy (ash2326):

Let's simplify the term inside the sine \[\sin (3(x+120))=\frac 12\] \[\sin (3x+360)=\frac 12\] We know that sine has a period of 360 degrees or \(\pi\) radians, so \[\sin (3x+360)=\sin 3x\] Now we get \[\sin (3x)=\frac 12\] Hey @bronzegoddess for what value of angle is sin \(\frac 12\)??

OpenStudy (anonymous):

30 or \[\Pi/6\]

OpenStudy (anonymous):

@ash2326 where did this come from?sin(3x+360)=sin3x

OpenStudy (ash2326):

@bronzegoddess it's the property of all the tangent functions, they repeat their value after a period of 360 or \(2\pi\). Sorry I missed the 2 earlier

OpenStudy (anonymous):

i know but then you just remove it from the equation?

OpenStudy (ash2326):

 o and 360 come at the same point, so if we have sin 390= sin 30 yeah, you can remove it just like that